Divide 4 1/5 by 3 1/7 Write the result in simplest form. I must show the work and I do not know how to reduce to simplest form.
Answer by edjones(8007)

4 1/3 / 3 1/7
=13/3 / 22/7
=13/3 * 7/22
=91/66
.
Ed